# Generative Auto-Bidding with Value-Guided Explorations

### v1

@@ Abstract.
Auto-bidding, with its strong capability to optimize bidding decisions within dynamic and competitive online environments, has become a pivotal strategy for advertising platforms. Existing approaches typically employ rule-based strategies or Reinforcement Learning (RL) techniques. However, rule-based strategies lack the flexibility to adapt to time-varying market conditions, and RL-based methods struggle to capture essential historical dependencies and observations within Markov Decision Process (MDP) frameworks. Furthermore, these approaches often face challenges in ensuring strategy adaptability across diverse advertising objectives. Additionally, as offline training methods are increasingly adopted to facilitate the deployment and maintenance of stable online strategies, the issues of documented behavioral patterns and behavioral collapse resulting from training on fixed offline datasets become increasingly significant. To address these limitations, this paper introduces a novel offline Generative Auto-bidding framework with Value-Guided Explorations (GAVE). GAVE accommodates various advertising objectives through a score-based Return-To-Go (RTG) module. Moreover, GAVE integrates an action exploration mechanism with an RTG-based evaluation method to explore novel actions while ensuring stability-preserving updates. A learnable value function is also designed to guide the direction of action exploration and mitigate Out-of-Distribution (OOD) problems. Experimental results on two offline datasets and real-world deployments demonstrate that GAVE outperforms state-of-the-art baselines in both offline evaluations and online A/B tests. The implementation code is publicly available to facilitate reproducibility and further research  [^1] https://anonymous.4open.science/r/GAVE-0404.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 48th ACM SIGIR Conference on Research and Development in Information Retrieval.; July 13–18, 2025; Padua, Italyisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Computational advertising
@@ 1. Introduction
Bidding remains a fundamental component of modern online advertising platforms (Zhao et al., 2018a, 2020; Gao et al., 2024b, a), enabling businesses to engage target audiences and increase sales. As digital advertising evolves, traditional manual bid management methods have become increasingly inadequate and cost-ineffective, failing to address the demands of today’s dynamic market (Jha et al., 2024). The complexity of modern advertising systems, characterized by fluctuating market conditions and diverse user behaviors (Aggarwal et al., 2024), requires automated solutions that can adapt to these variations while aligning with advertisers’ diverse objectives (Borissov et al., 2010; Wang et al., 2023a). This need is further intensified by the massive volume of ad impressions requiring real-time processing, where human intervention becomes impractical and often suboptimal for achieving advertising goals (Borissov et al., 2010; Guo et al., 2024).
@@
  • Experiments on two public datasets, along with results from online deployments, demonstrate the effectiveness of GAVE compared to various state-of-the-art offline bidding baselines.
@@ 3.2. Score-based RTG
Furthermore, in practical applications, different advertising objectives may exhibit varying degrees of dependence on CPA constraints, resulting in evaluation metrics that differ from Equation 13. Nonetheless, training and evaluation can remain aligned by generalizing the evaluation score $S$ to each time step (i.e., $S_{t}$) in a similar way, leading to:
@@ 3.3. Action Explorations
where $DT()$ represents the Decision Transformer backbone (a causal transformer as shown in Figure 1 (a)), $FC_{\beta}()$ denotes a fully-connected layer, and $\sigma$ is scaling function. To mitigate OOD issues, the scaling function is defined as:
@@ 3.4. Learnable Value Function
where $\mathbb{A}$ denotes the available action space. Due to the extensive action space inherent in bidding and the limited diversity of actions within offline datasets, the direct statistical computation of $V_{t+1}$ is infeasible. However, we can learn this value through an expectile regression process with $r_{t+1}$:
@@
By estimating $V_{t+1}$ with $\hat{V}_{t+1}$ and utilizing it to guide the updating directions of $\tilde{r}_{t+1}$, GAVE implicitly steers the update direction of the explored action $\tilde{a}_{t}$ toward potentially optimal actions for strategy improvement. This process is illustrated in Figure 1 (b.2) and can be formalized as:
(22) ${}L_{v}=\frac{1}{M+1}\sum_{t-M}^{t}{(\tilde{r}_{t+1}-\hat{V}^{\prime}_{t+1})^{2}}$
@@ 4.4. Parameter Analysis (RQ3)
From Figure 2, it is evident that as training progresses and the overall loss $L_{o}$ decreases, the parameter $w_{t}$ increases from approximately 0.5 to a stable position above 0.5. The stable position of $w_{t}$ is influenced by both the dataset distribution and the model’s hyper-parameters. This trend confirms the efficacy of the learnable value function in directing action exploration. With the guidance of the value function, the model consistently explores actions $\tilde{a}$ with higher RTG values $\tilde{r}_{t+1}$ near the estimated optimal value $\hat{V}_{t+1}$. This approach facilitates learning potential optimal strategies while mitigating OOD issues.
@@ 5. Online Application
For both Nobid and Costcap, GAVE improves cost and conversions. In Nobid campaigns, GAVE achieves a 0.414% increase in cost, 2.2% in conversions, and 4.4% in target cost. For Costcap campaigns, advertising revenue and advertiser value rise alongside a significant improvement in CPA validity, with +1.899% cost, +19.2% conversions, +2.08% target cost, and +5.76% valid CPA ratio.
@@ 6. Related Works
This section provides a brief review of relevant research topics, i.e., auto-bidding, and reinforcement learning.
@@ 6.2. Auto-Bidding at Online Advertising Platforms
Auto-bidding plays a vital role in managing large-scale ad auctions by automatically optimizing bids per impression to meet advertisers’ goals. Early methods like PID (Chen et al., 2011) and OnlineLP (Yu et al., 2017) focus on rule-based approaches, using feedback loops and stochastic programming to address budget pacing and bid optimization, though they depend on simplified assumptions. To tackle the complexity of modern advertising ecosystems, RL-based solutions such as RLB (Cai et al., 2017), USCB (He et al., 2021), MAAB (Wen et al., 2022), and SORL (Mou et al., 2022) enable adaptive decision-making with capabilities for handling high-dimensional states and multi-agent coordination. However, due to the risks of real-time bidding, offline RL methods like BCQ (Fujimoto et al., 2019), CQL (Kumar et al., 2020), and IQL (Kostrikov et al., [n. d.]) have gained prominence for leveraging historical data. These approaches, while effective under MDP frameworks, struggle with modeling long-range temporal dependencies crucial for sequential bid optimization. Generative sequential modeling methods such as DiffBid (Guo et al., 2024) and GAS (Li et al., 2024) address these challenges using diffusion models and MCTS with transformer critics for improved bid trajectory generation. In this work, we propose a score-based RTG, action exploration mechanisms, and a learnable value function framework to align optimization objectives with evaluation metrics, enhance action exploration, and learn optimal strategies for improving bidding performance.
@@ 7. Conclusion
In this work, we propose GAVE to enhance DTs for offline generative auto-bidding through value-guided explorations. To accommodate complex advertising objectives, we design a customizable score-based RTG mechanism, enabling adaptive modeling of diverse optimization objectives to align with different evaluation metrics. Moreover, we integrate an action exploration mechanism with an RTG-based evaluation method to explore actions outside the offline dataset while ensuring a stability-preserving update process. To further guide the exploration and mitigate OOD risks, we employ a learnable value function to anchor RTG updates to distributionally plausible regions while allowing controlled extrapolation for strategy improvement. Extensive experiments and online deployments demonstrate the effectiveness of our GAVE framework in enhancing the adaptability and performance of auto-bidding strategies, providing a versatile solution for optimizing digital advertising campaigns in dynamic environments.

### v2

@@ Abstract.
Auto-bidding, with its strong capability to optimize bidding decisions within dynamic and competitive online environments, has become a pivotal strategy for advertising platforms. Existing approaches typically employ rule-based strategies or Reinforcement Learning (RL) techniques. However, rule-based strategies lack the flexibility to adapt to time-varying market conditions, and RL-based methods struggle to capture essential historical dependencies and observations within Markov Decision Process (MDP) frameworks. Furthermore, these approaches often face challenges in ensuring strategy adaptability across diverse advertising objectives. Additionally, as offline training methods are increasingly adopted to facilitate the deployment and maintenance of stable online strategies, the issues of documented behavioral patterns and behavioral collapse resulting from training on fixed offline datasets become increasingly significant. To address these limitations, this paper introduces a novel offline Generative Auto-bidding framework with Value-Guided Explorations (GAVE). GAVE accommodates various advertising objectives through a score-based Return-To-Go (RTG) module. Moreover, GAVE integrates an action exploration mechanism with an RTG-based evaluation method to explore novel actions while ensuring stability-preserving updates. A learnable value function is also designed to guide the direction of action exploration and mitigate Out-of-Distribution (OOD) problems. Experimental results on two offline datasets and real-world deployments demonstrate that GAVE outperforms state-of-the-art baselines in both offline evaluations and online A/B tests. By applying the core methods of this framework, we proudly secured first place in the NeurIPS 2024 competition, ‘AIGB Track: Learning Auto-Bidding Agents with Generative Models’  [^1] https://tianchi.aliyun.com/competition/entrance/532236/rankingList. The implementation code is publicly available to facilitate reproducibility and further research  [^2] https://github.com/Applied-Machine-Learning-Lab/GAVE.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3729987isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Computational advertising
@@ 1. Introduction
Bidding remains a fundamental component of modern online advertising platforms (Zhao et al., 2018a, 2020; Gao et al., 2024b, a), enabling businesses to engage target audiences and increase sales. As digital advertising evolves, traditional manual bid management methods have become increasingly inadequate and cost-ineffective, failing to address the demands of today’s dynamic market (Jha et al., 2024; Liu et al., 2024). The complexity of modern advertising systems, characterized by fluctuating market conditions and diverse user behaviors (Aggarwal et al., 2024), requires automated solutions that can adapt to these variations while aligning with advertisers’ diverse objectives (Borissov et al., 2010; Wang et al., 2023a). This need is further intensified by the massive volume of ad impressions requiring real-time processing, where human intervention becomes impractical and often suboptimal for achieving advertising goals (Borissov et al., 2010; Guo et al., 2024).
@@
  • Experiments on two public datasets, along with results from online deployments, demonstrate the effectiveness of GAVE compared to various state-of-the-art offline bidding baselines. Additionally, by applying the core methods of this framework, we proudly secured first place in the NeurIPS 2024 competition, ‘AIGB Track: Learning Auto-Bidding Agents with Generative Models’ hosted by Alimama 11footnotemark: 1.
@@ 3.2. Score-based RTG
Furthermore, in practical applications, different advertising objectives may exhibit varying degrees of dependence on CPA constraints, resulting in different evaluation metrics. Nonetheless, training and evaluation can remain aligned by generalizing $S$ to each time step (i.e., $S_{t}$) in a similar way, leading to:
@@ 3.3. Action Explorations
where $DT()$ represents the DT backbone, $FC_{\beta}()$ denotes a fully-connected layer, and $\sigma$ is scaling function. To mitigate OOD issues, the scaling function is defined as:
@@ 3.4. Learnable Value Function
where $\mathbb{A}$ denotes the available action space. Due to the extensive action space and limited real actions within offline datasets, the direct statistical computation of $V_{t+1}$ is infeasible. However, we can learn this value through an expectile regression process with $r_{t+1}$:
@@
By estimating $V_{t+1}$ with $\hat{V}_{t+1}$ and utilizing it to guide the updating directions of $\tilde{r}_{t+1}$, GAVE implicitly steers the update direction of the explored $\tilde{a}_{t}$ toward potentially optimal actions. This process is illustrated in Figure 1 (b.2) and can be formalized as:
(22) ${}L_{v}=\frac{1}{M+1}\sum_{t-M}^{t}{(\tilde{r}_{t+1}-\hat{V}^{\prime}_{t+1})^{2}}$
@@ 4.4. Parameter Analysis (RQ3)
From Figure 2, it is evident that as training progresses, the parameter $w_{t}$ increases from approximately 0.5 to a stable position above 0.5. The stable position is influenced by both the dataset distribution and the model’s hyper-parameters. This trend confirms the efficacy of the learnable value function in directing action exploration. With the guidance of the value function, the model consistently explores actions $\tilde{a}$ with higher RTG values $\tilde{r}_{t+1}$ near the estimated optimal value $\hat{V}_{t+1}$. This approach facilitates learning potential optimal strategies while mitigating OOD issues.
@@ 5. Online Application
For both Nobid and Costcap, GAVE improves cost and conversions. In Nobid campaigns, GAVE achieves a 0.8% increase in cost, 8.0% in conversions, and 3.2% in target cost. For Costcap campaigns, advertising revenue and advertiser value rise alongside a significant improvement in CPA validity, with +2.0% cost, +3.6% conversions, +2.2% target cost, and +1.9% valid CPA ratio.
@@ 6. Related Works
This section provides a brief review of relevant research topics, i.e., offline reinforcement learning, and auto-bidding.
@@ 6.2. Auto-Bidding at Online Advertising Platforms
Auto-bidding plays a vital role in managing large-scale ad auctions by automatically optimizing bids per impression to meet advertisers’ goals. Early methods like PID (Chen et al., 2011) and OnlineLP (Yu et al., 2017) focus on rule-based approaches, using feedback loops and stochastic programming to address budget pacing and bid optimization, though they depend on simplified assumptions. To tackle the complexity of modern advertising ecosystems, RL-based solutions such as RLB (Cai et al., 2017), USCB (He et al., 2021), MAAB (Wen et al., 2022), and SORL (Mou et al., 2022) enable adaptive decision-making with capabilities for handling high-dimensional states and multi-agent coordination. However, due to the risks of real-time bidding, offline RL methods like BCQ (Fujimoto et al., 2019), CQL (Kumar et al., 2020), and IQL (Kostrikov et al., [n. d.]) have gained prominence for leveraging historical data. These approaches, while effective under MDP frameworks, struggle with modeling long-range temporal dependencies crucial for sequential bid optimization. Generative sequential modeling methods such as DiffBid (Guo et al., 2024) and GAS (Li et al., 2024) address these challenges using diffusion models and transformers with MCTS for improved bid trajectory generation. In this work, we propose a score-based RTG, action exploration mechanisms, and a learnable value function framework to align optimization objectives with evaluation metrics, enhance action exploration, and learn optimal strategies for improving bidding performance with Decision Transformer.
@@ 7. Conclusion
In this work, we propose GAVE to enhance DTs for offline generative auto-bidding through value-guided explorations. To accommodate complex advertising objectives, we design a customizable score-based RTG mechanism, enabling adaptive modeling of diverse optimization objectives to align with different evaluation metrics. Moreover, we integrate an action exploration mechanism with an RTG-based evaluation method to explore actions outside the offline dataset while ensuring a stability-preserving update process. To further guide the exploration and mitigate OOD risks, we employ a learnable value function to anchor RTG updates to distributionally plausible regions while allowing controlled extrapolation for strategy improvement. Extensive experiments, online deployments and NeurIPS competition 11footnotemark: 1 demonstrate the effectiveness of our GAVE framework in enhancing the adaptability and performance of auto-bidding strategies, providing a versatile solution for optimizing digital advertising campaigns in dynamic environments.